OpenStudy (ashsmashum):

Determaine whether each equation represents a direct variation if it does find the constant variation 2y=5x+1

OpenStudy (mathmale):

Solve 2y=5x+1 for y, so that you can tell more quickly whether or not this is an example of direct variation. Divide both sides of this equation by 2. Your result?

OpenStudy (ashsmashum):

\[2y \div 2\]

OpenStudy (ashsmashum):

?

OpenStudy (ashsmashum):

I'm sorry if I'm wrong i'm pretty bad at math

OpenStudy (mathmale):

\[\frac{ 2y }{ 2 }=\frac{ 5x }{ 2 }+\frac{ 1 }{ 2 }=?\]

OpenStudy (ashsmashum):

2y/2=1

OpenStudy (ashsmashum):

i keep on getting 2/5

OpenStudy (ashsmashum):

for the second 5/2

OpenStudy (mathmale):

Note: (2y)/2 is not 1; it is y. 2y -- = y 2 So you have y=(5/2)x + (1/2). Example of DIRECT PROPORTION: y=8x Example of first order function that is NOT direct proportion: y = mx + b, or y=6x - 7. So, is the given function an example of direct proportion, or NOT an example?

OpenStudy (mathmale):

Please, try to complete your work on problems you post. When done, please close your post and thank the person or persons who have helped you. Thanks.

OpenStudy (ashsmashum):

sorry i went offline